OpenStudy (yanasidlinskiy):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (yanasidlinskiy):

This is what I have: 

OpenStudy (anonymous):

the lateral área is like área of a rectangle with hight=50 and length equal to circle length = 2\(\pi\)r are of base is área of circle = \(\pi r^2\), so: adding this two you need to get around 3000: (2\(\pi r\cdot \)50)+\(\pi r^2\)=3000 this a quadratic equation. Solve it for r

OpenStudy (mathmate):

Area of the base (assuming a vase is open at the top) \[=\pi r^2\ cm^2\] Lateral Area \[= 2\pi rh\ cm^2\] \[= 100\pi r\ cm^2\] If the total surface area is 3000 cm2, then \[\pi r^2 + 100\pi r = 3000\] Form the following quadratic equation and solve for r: \[\pi r^2 + 100\pi r - 3000 = 0\] to give r=8.77 cm or -108... Reject the illogical value.